# Congruent modular forms and anticyclotomic Iwasawa theory

### v1

@@ 5.2 Hecke operators at $\ell\neq p$ in Serre-Tate coordinates
For $\gcd(N,\ell)\mid N^{\sharp}$, we may naturally identify $f$ as a form of level $N^{\sharp}$. For an ordinary test triplet $(A,\eta_{N^{\sharp}},\omega)\in\mathcal{M}_{\Gamma_{1}(N^{\sharp})}$ with level $N^{\sharp}$ structure $\eta_{N^{\sharp}}$, let $C\subset A[N^{\sharp}]$ be the image of the level structure $\eta_{N^{\sharp}}$.
@@
has kernel $\mu_{\ell}$, and we will denote by $\overline{\pi\circ\eta_{N^{\sharp}}}$ the isomorphism
@@
Moreover, we will denote by $(\cdot)^{1/\ell}$ the inverse of the isomorphism $\mu_{N^{\sharp}}/\mu_{\ell}\xrightarrow{\zeta\rightarrow\zeta^{\ell}}\mu_{N^{\sharp}\ell^{-1}}$. One can define the ’dividing by $\ell$-level structure’ operator $V_{\ell}$ on ordinary test triplets as
@@
where $\pi:A\rightarrow A/C[\ell]$ is the canonical projection and $\check{\pi}:A/C[\ell]\rightarrow A$ is its dual isogeny.
This induces an operator $V_{\ell}^{*}$ on the space of classical modular forms of level $\Gamma_{0}(N^{\sharp})$ via the rule $V_{\ell}^{*}f(A,\eta_{N^{\sharp}},\omega)=f(V_{\ell}(A,\eta_{N^{\sharp}},\omega))$, which acts on $q$-expansions as $f(q)\mapsto f(q^{\ell})$ [10, 14, 15].
@@
We define the $(\ell)$-stabilization for a newform $f$ of conductor $N$ and weight $2r$ as:
@@
where we implicitly identify $f$ as a form of level $N^{\sharp}$.
@@
$[N^{\sharp}/N]^{\ast}:V_{p}(N,\mathcal{W})\hookrightarrow V_{p}(N^{\sharp},\mathcal{W}).$
For $\gcd(N,\ell)\mid N^{\sharp}$, we define the following analogue of the $V_{\ell}$ operator for $p$-adic modular forms:
@@
For a complete local $\mathcal{W}$-algebra $R$ and $[(A,\eta)_{/R}]\in\text{Ig}(N)(R)$, recall from Section 2 that the $\Gamma_{1}(Np^{\infty})$-level structure $\eta=\eta_{N}\oplus\eta_{p}$ determines a map $\widehat{\eta}_{p}:\widehat{\mathbf{G}}_{m}\xrightarrow{\sim}\widehat{A}$ [14, (1.10.1)](see also [15, Proposition 1]), which defines a differential $\omega(\widehat{\eta}_{p}):\text{Lie}(A)\simeq\text{Lie}(\widehat{A})\rightarrow\text{Lie}(\widehat{\mathbf{G}}_{m})=R$. A geometric modular form $f$ can then be identified as a $p$-adic modular form via the rule $\widehat{f}(A,\eta)=f(A,\eta,\omega(\widehat{\eta_{p}}))$. To show the compatibility of the $V_{\ell}$ operator defined on geometric modular forms and $p$-adic modular forms, we begin with the following
@@
Let $\phi:A/R\longrightarrow A^{\prime}/R$ be an isogeny of elliptic curves. Suppose that $\eta_{p}$ is a $p^{\infty}$-level structure on $A/R$, and $\phi\circ\eta_{p}$ is the $p^{\infty}$- level structure on $A^{\prime}/R$ induced by $\phi$. Then $(\phi^{\ast})^{-1}\omega(\widehat{\eta_{p}})=\omega(\widehat{\phi\circ\eta_{p}})$, where $\phi^{\ast}:H^{0}(A^{\prime}/R,\underline{\Omega}_{A^{\prime}/R}^{1})\rightarrow H^{0}(A/R,\underline{\Omega}_{A/R}^{1})$ is induced by $\phi$.
@@
Recall that $\omega(\widehat{\eta_{p}})$ (respectively $\omega(\widehat{\phi\circ\eta_{p}})$) is defined as the inverse of $\text{Lie}(\widehat{\eta}_{p})$ (respectively $\text{Lie}(\widehat{\phi\circ\eta_{p}})$):
@@
Hence, we have $\text{Lie}(\widehat{\phi})^{\ast}\omega(\widehat{\phi\circ\eta_{p}})=\omega(\widehat{\eta}_{p})$ by functoriality, where $\text{Lie}(\widehat{\phi})^{\ast}$ is the pull-back map induced by $\text{Lie}(\widehat{\phi})$. Moreover, the map $\text{Lie}(\widehat{\phi})^{\ast}$ is the same as $\phi^{\ast}:H^{0}(A^{\prime}/R,\underline{\Omega}_{A^{\prime}/R}^{1})\rightarrow H^{0}(A/R,\underline{\Omega}_{A/R}^{1})$, and we have $\phi^{\ast}\omega(\widehat{\phi\circ\eta_{p}})=\omega(\widehat{\eta}_{p}).$
@@
where the $V_{\ell}^{\ast}$ operator on the left-hand side acts on $p$-adic modular forms and the $V_{\ell}^{\ast}$ operator on the right-hand side acts on geometric modular forms.
@@
This follows directly from $\check{\pi}^{\ast}\omega(\widehat{\eta_{p}})=\omega(\widehat{\check{\pi}^{-1}\circ\eta_{p}})$ by the previous lemma, and the definitions
@@
For ease of notation, we will denote $A=A_{\mathfrak{a}}$ and $\eta=\eta_{\mathfrak{a}}$. We denote by $\pi_{v}$ the projection $A\rightarrow A/A[v]$, and by $\pi_{\overline{v}}$ the projection $A_{v}\rightarrow A_{v}/A_{v}[\overline{v}]$ where $A_{v}=A/A[v]$. Observe that
@@
introduced in [10, Lemma 3.5] gives rise to the isomorphism between the tuples
@@
Indeed, following the argument of [10, Lemma 3.5], the composition $\iota\circ\pi_{\overline{v}}\circ\pi_{v}$ is the multiplication by $\ell$ map $[\ell]:A\rightarrow A$. This implies that the dual isogeny $\check{\pi}_{v}$ of $\pi_{v}$ is $\iota\circ{{\pi}_{\overline{v}}}$, so that $\iota\circ\pi_{\overline{v}}\circ\check{\pi}_{v}^{-1}\circ\eta_{p}=\eta_{p}.$
@@
Let $\phi:\mathbb{Z}_{p}^{\times}\rightarrow\mathcal{O}_{\mathbb{C}_{p}^{\times}}$ be a $p$-adic character of conductor $p^{n}$. We have the following identity: $((\theta^{-r}V_{\ell}^{\ast}f)_{\mathfrak{a}}\otimes\phi)(x_{\mathfrak{a}})=\ell^{-r}((\theta^{-r}f)_{\overline{v}^{-1}\mathfrak{a}}\otimes\phi)(x_{\overline{v}^{-1}\mathfrak{a}}).$
@@
Following [1, 8, 9], we define $\mathscr{P}_{v}\in\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket$ such that
@@
where $\gamma_{v}\in\Gamma_{K}^{-}$ is the Frobenius at $v$. We define $\mathscr{P}_{\overline{v}}(f)\in\mathcal{W}\llbracket\tilde{\Gamma}\rrbracket$ similarly.
We fix a topological generator $\gamma_{0}$ of $\Gamma_{K}^{-}$, and let $\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket\simeq\mathcal{W}\llbracket T\rrbracket$ be the isomorphism given by $\gamma_{0}\mapsto T+1$.
@@
As elements in $\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket$, both $\mathscr{P}_{v}(f)$ and $\mathscr{P}_{\overline{v}}(f)$ have $\mu$-invariants 00.
@@
We can similarly show that $\mu(\mathscr{P}_{\ell}(f))=0$ for $\ell\nmid N$. Indeed, we may write $\mathscr{P}_{v}(f)=(1-a_{\ell}\cdot\gamma_{v})(1-b_{\ell}\cdot\gamma_{v})$, and it can be shown by the same argument as above that both $1-a_{\ell}\cdot\gamma_{v},1-b_{\ell}\cdot\gamma_{v}$ have $\mu$-invariants 00. The same argument applies to $\mathscr{P}_{\overline{v}}(f)$. ∎
@@
For every locally algebraic character $\rho\in\mathfrak{X}_{p^{\infty}}$, we use Lemma 5.6 to obtain the following:
@@
Suppose that $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ are newforms satisfying Hypothesis (1) whose coefficients lie in some number field $L$. We assume that $\mathcal{W}$ is the ring of integers of a finite extension of $\widehat{\mathbb{Q}_{p}^{nr}}$ containing $L$.
Suppose that the induced semi-simplified $\mod{\varpi}^{m}$ Galois representations: $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:\operatorname{Gal}(\overline{L}/L)\rightarrow GL_{2}(\mathcal{O}_{L_{p}}/\varpi^{m}\mathcal{O}_{L_{p}})$ are isomorphic, where $\varpi$ is the uniformizer of $\mathcal{O}_{L_{p}}$. For each prime $\ell\mid N_{1}N_{2}$, let $v\mid\mathfrak{N}_{1}\mathfrak{N}_{2}$ be the corresponding prime above $\ell$. Then the following congruence holds:
@@
Note that for $F\in\mathcal{W}\llbracket\Gamma_{K}^{-}\llbracket\simeq\mathcal{W}\rrbracket T\rrbracket$ with $\mu(F)=0$, we have $\lambda(F)=\lambda(\overline{F})$, where $\overline{F}\in\bar{\mathbb{F}}_{p}\llbracket T\rrbracket$ is the reduction of $F$ $\mod{\varpi}$. When $\mu(\mathscr{L}_{\mathfrak{p}}(f_{1}))=\mu(\mathscr{L}_{\mathfrak{p}}(f_{2}))=0$, it follows that
@@ 6 Applications to generalized Heegner cycles
6.1 Definitions
In this section, we follow the set-up of [5, Section 4]. As before, let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a normalized Hecke eigenform of weight $2r$ and level $N$ satisfying hypothesis (1).
We may write $K=\mathbb{Q}(\sqrt{-D_{K}})$ where $D_{K}$ is the discriminant of $K$, and for $r>1$ assume that either $-D_{K}>3$ is odd, or $8\mid D_{K}$. Such an assumption gurantees a canonical choice of elliptic curve $A$ with CM by $\mathcal{O}_{K}$, defined over the real subfield of the Hilbert class field $H_{K}$ of $K$ [5, Section 4.1].
Recall that $V=V_{f}(r)$ is the self-dual $p$-adic Galois representation associated with $f$. For primes $p$ such that $p\nmid 2(2r-1)!N\phi(N)$, we denote by $z_{f,\chi}\in H^{1}_{f}(K,T\otimes\chi)$ the generalized Heegner class attached to $(f,\chi)$ as constructed in [5, Section 4.5]. We remark that the construction involves the aforementioned canonical CM elliptic curve $A$.
We recall the definition of the Bloch-Kato logarithm map. Let $\mathbf{B}_{\text{dR}},\mathbf{B}_{\text{cris}}$ be Fontaine’s rings of $p$-adic periods [21, Definition 5.15, Definition 6.7], and let $t\in\mathbf{B}_{\text{dR}}$ be Fontaine’s $p$-adic analogue of $2\pi i$ [21, Section 5.2.3].
Denote by $\mathbf{D}_{\text{dR},L}(V)$ the filtered $(L\otimes_{\mathbb{Q}_{p}}F)$-module $(\mathbf{B}_{\text{dR}}\otimes V)^{G_{L}}$ and define $H^{1}_{f}(L,V):=\ker(H^{1}(L,V)\rightarrow H^{1}(L,\mathbf{B}_{\text{cris}}\otimes V))$ in accordance with [22, (3.7.2)]. The following exponential map is due to Bloch and Kato [22, Section 3]:

### v2

@@ 5.2 Hecke operators at $\ell\neq p$ in Serre-Tate coordinates
Let $\ell\neq p$ be a rational prime. For $\operatorname{lcm}(N,\ell)\mid N^{\sharp}$, one may naturally identify $f$ as a form of level $N^{\sharp}$. For an ordinary test triplet $(A,\eta_{N^{\sharp}},\omega)\in\mathcal{M}_{\Gamma_{1}(N^{\sharp})}$ with level $N^{\sharp}$ structure $\eta_{N^{\sharp}}$, let $C\subset A[N^{\sharp}]$ be the image of the level structure $\eta_{N^{\sharp}}$.
@@
has kernel $\mu_{\ell}$, and denote by $\overline{\pi\circ\eta_{N^{\sharp}}}$ the isomorphism
@@
Moreover, denote by $(\cdot)^{1/\ell}$ the inverse of the isomorphism $\mu_{N^{\sharp}}/\mu_{\ell}\xrightarrow{\zeta\rightarrow\zeta^{\ell}}\mu_{N^{\sharp}\ell^{-1}}$.
Definition 5.2.
Define the ’dividing by $\ell$-level structure’ operator $V_{\ell}$ on ordinary test triplets as
@@
where $\pi:A\rightarrow A/C[\ell]$ is the canonical projection and $\check{\pi}:A/C[\ell]\rightarrow A$ is its dual isogeny. Here, $\check{\pi}^{*}$ is the map on differential forms induced by $\check{\pi}$.
The map $V_{\ell}$ induces an operator $V_{\ell}^{*}$ on the space of classical modular forms of level $\Gamma_{0}(N^{\sharp})$ via the rule $V_{\ell}^{*}f(A,\eta_{N^{\sharp}},\omega)=f(V_{\ell}(A,\eta_{N^{\sharp}},\omega))$, which acts on $q$-expansions as $f(q)\mapsto f(q^{\ell})$ [10, 14, 15].
@@
Define the $(\ell)$-stabilization for a newform $f$ of conductor $N$ and weight $2r$ as:
@@
where $f$ is viewed as a form of level $N^{\sharp}$.
@@
$N^{\sharp}/N]^{\ast}:V_{p}(N,\mathcal{W})\hookrightarrow V_{p}(N^{\sharp},\mathcal{W}).$
Definition 5.4.
For $\operatorname{lcm}(N,\ell)\mid N^{\sharp}$, define the following analogue of the $V_{\ell}$ operator for $p$-adic modular forms:
@@
For a complete local $\mathcal{W}$-algebra $R$ and $(A,\eta)_{/R}]\in\text{Ig}(N)(R)$, recall from Section 2 that the $\Gamma_{1}(Np^{\infty})$-level structure $\eta=\eta_{N}\oplus\eta_{p}$ determines a map $\widehat{\eta}_{p}:\widehat{\mathbf{G}}_{m}\xrightarrow{\sim}\widehat{A}$ [14, (1.10.1)](see also [15, Proposition 1]), which defines a differential $\omega(\widehat{\eta}_{p}):\text{Lie}(A)\simeq\text{Lie}(\widehat{A})\rightarrow\text{Lie}(\widehat{\mathbf{G}}_{m})=R$. A geometric modular form $f$ can then be identified as a $p$-adic modular form via the rule $)\widehat{f}(A,\eta)=f(A,\eta,\omega(\widehat{\eta_{p}}$)). To show the compatibility of the $V_{\ell}$ operator defined on geometric modular forms and $p$-adic modular forms, we begin with the following
@@
[…]
@@
[…]
@@
[…]
@@
where the $V_{\ell}^{\ast}$ operator on the left-hand side acts on $p$-adic modular forms (Definition 5.4) and the $V_{\ell}^{\ast}$ operator on the right-hand side acts on geometric modular forms (Definition 5.2).
@@
[…]
@@
For ease of notation, we will denote $A=A_{\mathfrak{a}}$ and $\eta=\eta_{\mathfrak{a}}$ in this proof. Denote by $\pi_{v}$ the projection $A\rightarrow A/A[v]$, and by $\pi_{\overline{v}}$ the projection $A_{v}\rightarrow A_{v}/A_{v}[\overline{v}]$ where $A_{v}=A/A[v]$. Observe that
@@
introduced in Lemma 3.5 of [10] gives rise to the isomorphism between the tuples
@@
Indeed, following the argument in loc. cit., the composition $\iota\circ\pi_{\overline{v}}\circ\pi_{v}$ is the multiplication by $\ell$ map $\ell]:A\rightarrow A$. This implies that the dual isogeny $\check{\pi}_{v}$ of $\pi_{v}$ is $\iota\circ{{\pi}_{\overline{v}}}$, so that $\iota\circ\pi_{\overline{v}}\circ\check{\pi}_{v}^{-1}\circ\eta_{p}=\eta_{p}.$
@@
Let $\phi:\mathbb{Z}_{p}^{\times}\rightarrow\mathcal{O}_{\mathbb{C}_{p}^{\times}}$ be a $p$-adic character of conductor $p^{n}$. We have the following identity:
$((\theta^{-r}V_{\ell}^{\ast}f)_{\mathfrak{a}}\otimes\phi)(x_{\mathfrak{a}})=\ell^{-r}((\theta^{-r}f)_{\overline{v}^{-1}\mathfrak{a}}\otimes\phi)(x_{\overline{v}^{-1}\mathfrak{a}}).$
@@
Following [1, 8, 9], define $\mathscr{P}_{v}\in\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket$ such that
@@
where $\gamma_{v}\in\Gamma_{K}^{-}$ is the Frobenius at $v$. Define $\mathscr{P}_{\overline{v}}(f)\in\mathcal{W}\llbracket\tilde{\Gamma}\rrbracket$ similarly.
Fix a topological generator $\gamma_{0}$ of $\Gamma_{K}^{-}$, and let $\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket\simeq\mathcal{W}\llbracket T\rrbracket$ be the isomorphism given by $\gamma_{0}\mapsto T+1$.
@@
As elements in $\mathcal{W}\llbracket\Gamma_{K}^{-}\rrbracket$, both $\mathscr{P}_{v}(f)$ and $\mathscr{P}_{\overline{v}}(f)$ have $\mu$-invariants 0.
@@
We can similarly show that $\mu(\mathscr{P}_{\ell}(f))=0$ for $\ell\nmid N$. Indeed, we may write $\mathscr{P}_{v}(f)=(1-a_{\ell}\cdot\gamma_{v})(1-b_{\ell}\cdot\gamma_{v})$, and it can be shown by the same argument as above that both $1-a_{\ell}\cdot\gamma_{v},1-b_{\ell}\cdot\gamma_{v}$ have $\mu$-invariants 0. The same argument applies to $\mathscr{P}_{\overline{v}}(f)$. ∎
@@
For every locally algebraic character $\rho\in\mathfrak{X}_{p^{\infty}}$, we use Lemma 5.8 to obtain the following identity:
@@
Suppose that $f_{1}\in S_{2r_{1}}(\Gamma_{0}(N_{1}))^{\text{new}},f_{2}\in S_{2r_{2}}(\Gamma_{0}(N_{2}))^{\text{new}}$ are newforms satisfying Hypothesis (1) whose coefficients lie in some $p$-adic field $F$. Assume that $\mathcal{W}$ is the ring of integers of a finite extension of $\widehat{\mathbb{Q}_{p}^{nr}}\cdot F$.
Suppose that the induced semi-simplified $\mod{\varpi}^{m}$ Galois representations: $\bar{\rho}_{f_{1}},\bar{\rho}_{f_{2}}:G_{\mathbb{Q}}\rightarrow\text{GL}_{2}(\mathcal{O}_{F}/\varpi^{m}\mathcal{O}_{F})$ are isomorphic, where $\varpi$ is the uniformizer of $\mathcal{O}_{F}$. For each prime $\ell\mid N_{1}N_{2}$, let $v\mid\mathfrak{N}_{1}\mathfrak{N}_{2}$ be the corresponding prime above $\ell$. Then the following congruence holds:
@@
Note that for an element $\mathscr{P}\in\mathcal{W}\llbracket\Gamma_{K}^{-}\llbracket\simeq\mathcal{W}\rrbracket T\rrbracket$ with $\mu(\mathscr{P})=0$, we have $\lambda(\mathscr{P})=\deg(\overline{\mathscr{P}})$, where $\overline{\mathscr{P}}\in\bar{\mathbb{F}}_{p}\llbracket T\rrbracket$ is the reduction of $\mathscr{P}$ $\mod{\varpi}$. When $\mu(\mathscr{L}_{\mathfrak{p}}(f_{1}))=\mu(\mathscr{L}_{\mathfrak{p}}(f_{2}))=0$, it follows that
@@ 6 Applications to generalized Heegner cycles
In this section, we follow the set-up of [3, Section 4]. As before, let $f\in S_{2r}(\Gamma_{0}(N))^{\text{new}}$ be a normalized Hecke eigenform of weight $2r$ and level $N$ satisfying hypothesis (1).
6.1 Generalized Heegner classes
Recall that $K=\mathbb{Q}(\sqrt{-D_{K}})$ where $D_{K}$ is the discriminant of $K$, and for $r>1$ assume that either $D_{K}>3$ is odd, or $8\mid D_{K}$. Such an assumption gurantees a canonical choice of elliptic curve $A$ with CM by $\mathcal{O}_{K}$, defined over the real subfield of the Hilbert class field $H_{K}$ of $K$ [3, Section 4.1].
Recall that $V=V_{f}(r)$ is the self-dual $p$-adic Galois representation associated with $f$. Let $T$ be a $G_{\mathbb{Q}}$-stable lattice in $V$. For primes $p$ such that $p\nmid 2(2r-1)!N\phi(N)$, denote by $z_{f,\chi}\in H^{1}_{f}(K,T\otimes\chi)$ the generalized Heegner class attached to $(f,\chi)$ [3, Section 4.5]. We remark that the construction involves the aforementioned canonical CM elliptic curve $A$.
6.2 Bloch-Kato logarithm map for a $p$-adic Galois representation
Now, we recall the definition of the Bloch-Kato logarithm map. Let $\mathbf{B}_{\text{dR}},\mathbf{B}_{\text{cris}}$ be Fontaine’s rings of $p$-adic periods [21, Definition 5.15, Definition 6.7], and let $t\in\mathbf{B}_{\text{dR}}$ be Fontaine’s $p$-adic analogue of $2\pi i$ [21, Section 5.2.3].
Let $F/\mathbb{Q}_{p}$ be a finite extension. Suppose that $V$ is an $F$-vectorspace that is also a $G_{L}$-module for some finite extension $L/\mathbb{Q}_{p}$ (such as the $\operatorname{Gal}(\overline{\mathbb{Q}_{p}}/\mathbb{Q}_{p})$-representation $V=V_{f}(r)$ where $F$ is a $p$-adic field containing the coefficients of $f$).
Denote by $\mathbf{D}_{\text{dR},L}(V)$ the filtered $(L\otimes_{\mathbb{Q}_{p}}F)$-module $(\mathbf{B}_{\text{dR}}\otimes V)^{G_{L}}$ and define $H^{1}_{f}(L,V):=\ker(H^{1}(L,V)\rightarrow H^{1}(L,\mathbf{B}_{\text{cris}}\otimes V))$ in accordance with [22, (3.7.2)]. If $V$ is a de Rham representation, the following exponential map is due to Bloch and Kato [22, Section 3]: